# MPCM-Net: Multi-scale network integrates partial attention convolution with Mamba for ground-based cloud image segmentation

### v1

@@ III-E Loss Function
Ground-based cloud image segmentation is fundamentally a binary pixel-level classification task. However, imbalanced data class distribution is a basic challenge in this task due to significant scale variations among cloud formations and substantial contrast disparities between certain cloud regions and the background, which biases models toward background identification. To mitigate this imbalance and accelerate network convergence during training, we employ a compound loss function, mathematically expressed as:
$L_{joint}=\alpha L_{focal}+\beta L_{dice}$ (25)
where $L_{focal}$ denotes the focal loss and $L_{dice}$ denotes the dice loss. $\alpha$ and $\beta$ are both hyperparameters, set to 0.6 and 0.4, respectively.
Focal loss addresses the limitation of class weight distribution in cross-entropy loss functions and is widely adopted for class-imbalanced tasks. By incorporating a modulating factor that assigns higher weights to hard-to-classify samples, it enhances the model’s focus on challenging cloud boundary details during training, particularly in ground-based cloud imagery segmentation. The specific formulation of the Focal loss is as follows:
$\begin{split}L_{focal}=-\frac{1}{N}\sum_{i}^{N}(\gamma(1-p_{i})^{\delta}\times p_{i}log(p_{i})+\\ (1-\gamma)p_{i}^{\delta}\times(1-g_{i})log(1-p_{i}))\end{split}$ (26)
where $p_{i}$ and $g_{i}$ denote the prediction and ground truth of sample $i$, respectively. $N$ denotes the total pixels of the input image. $\gamma$ and $\delta$ denote the weight of the positive and negative samples. They are hyperparameters and the values are set to 0.25 and 0.2, respectively, as same as [LG-Umer].
Dice loss mitigates foreground-background class imbalance by leveraging the Dice coefficient, where the model iteratively optimizes the similarity between predicted and ground-truth segmentation masks during training to balance cloud-background distributions. Given the significant scale variations and indistinct boundaries characteristic of cloud formations in ground-based cloud imagery, this loss function enhances foreground feature extraction, thereby increasing model sensitivity to critical cloud characteristics. The specific formulation of the Dice loss is as follows:
$L_{dice}=1-\frac{2\times\sum_{i}^{N}p_{i}g_{i}+\varepsilon}{\sum_{i}^{N}(p_{i}^{2}+g_{i}^{2})+\varepsilon}$ (27)
where $\varepsilon$ denotes the smoothing coefficient, which is set to $10^{-5}$ as a hyperparameter to prevent the denominator in equation (27) from being zero.
@@ IV-A Dataset
To address this gap and provide the research community with a more challenging and practical benchmark, we developed and publicly released a new dataset incorporating Complex-Scale variations, Radiative properties, and Color attributes (CSRC), which constitutes a core contribution of this paper. The solar radiation effect on clouds is a critical factor in the accuracy of PV power forecasting. International cloud classification standards are primarily based on the brightness, color, size, and altitude of clouds. Among these, color variations induced by radiative sources play a significant role in the precision of pixel-level classification in ground-based cloud imagery. In this study, we focus particularly on cloud scale and color attributes, and introduce fine-grained segmentation that incorporates radiative influence, re-categorizing clouds into three classes (see Table I for details). White and gray clouds are common color patterns in the sky, often covering extensive areas, and are represented by white and gray labels, respectively. Pixels near the sun often appear nearly pure white due to high radiation, with red and blue channel values converging, leading these regions to be classified as overcast. Therefore, it is essential to incorporate fine-grained solar radiation segmentation [dataset7], which is indicated in red. Black cloud systems, typically associated with heavy rain or hail, result in full sky coverage and are detrimental to both ground-based cloud image segmentation and PV power prediction tasks. Such black clouds are ignored from our dataset. Cloud-free sky regions are labeled in blue as background. A comparative analysis between our proposed CSRC dataset and existing public datasets is presented in the Table II, with evaluation criteria encompassing fine-grained categorization support, inclusion of radiation sources, color attribute availability, and spatial resolution.
@@
Dataset Fine-grained Presence of Radiation Source Color Attributes Resolution
@@
SWINSEG[dataset2] No No No $\times500$
WAHRSIS[dataset3] No No No $\times600$
Fabel et al.[dataset4] No No No $\times512$
Park et al.[dataset5] No No No $\times300$
TLCDD[dataset6] No No No $\times512$
@@
Shi et al.[CloudFU-Net] Yes No No $\times1704$
CSRC (Ours) Yes Yes Yes $\times1260$
The images in our CSRC dataset were collected at a meteorological observation station located in Xiqing District, Tianjin, China (geographic coordinates: 117.03° E, 39.10° N). The core acquisition device is an All-Sky Imager (ASI-DC-TK02), which provides a wide-angle field of view exceeding 180° and is housed in a waterproof and dustproof enclosure. The system was configured to automatically capture images at fixed 30-second intervals. All images are stored in RGB color JPG format with a resolution of 1260 × 1260 pixels. Representative samples from the CSRC dataset are shown in Fig. 7.
@@ IV-B Evaluation Metrics
Ground-based cloud image segmentation fundamentally constitutes a downstream task of semantic segmentation in computer vision. Consequently, to evaluate the effectiveness of our proposed method, we adopt common pixel-wise evaluation metrics, including Precision (P), Recall (R), and Mean Intersection over Union (MIoU). All the metrics are derived from true positives (TP), false positives (FP), and false negatives (FN) in experimental results, as specified in the following equations:
@@
where $p$ and $g$ denote the prediction results of our method and the ground truth, respectively, $N$ represents the number of categories in our dataset, and $i=\left\{0,1,2,3\right\}$ represents a total of four categories.
@@ IV-C Implementation Details
The proposed method was implemented in Python 3.9 using PyTorch 1.10.1 on an Ubuntu 21.04 server equipped with an Intel (R) Xeon (R) Gold 5218N CPU @ 2.30 GHz and an NVIDIA GeForce RTX 4090 GPU (24GB VRAM). Training employed an initial learning rate of $10^{-2}$ with the Adam optimizer, configured with momentum 0.9 and weight decay $10^{-3}$ to ensure stable convergence. All input images were uniformly cropped to 512 $\times$ 512 resolution, consistent with the literature [CloudFU-Net], to maintain experimental validity. Learning rate decay was applied every 10 epochs, with training termination triggered if no validation loss improvement occurred within 20 consecutive epochs. To enhance model generalization through input diversity, data augmentation strategies including random rotation (0°, 90°, 180°, 270°) and horizontal flipping (probability 0.3) were implemented. The complete training regimen required 8.9 hours for 50 epochs on our dataset.
@@ IV-D Results and Discussions
To evaluate the performance of the proposed MPCM-Net, we conduct comparative experiments against state-of-the-art methods for ground-based cloud segmentation. These methods include both classical semantic segmentation architectures (U-Net [U-Net], DeepLabv3+ [DeepLabv3], SegFormer [SegFormer]) and recent specialized approaches (CloudFU-Net [CloudFU-Net], CloudSwinNet [CloudSwinNet], MFAFNet [MFAFNet], BSANet [BSANet], and FA-CloudSeg [FA-CloudSeg]). All comparative results were obtained from publicly available implementations or original author-provided results, with experimental configurations rigorously aligned across methods to ensure fair evaluation.
@@ IV-D1 Quantitative Comparison
TABLE III: Quantitative Comparison with Different Methods on the CSRC Dataset. $\uparrow$ Indicates Higher is Better. The Best Results are Highlighted in Bold.
@@
U-Net (15MICCAI)[U-Net] 66.7 64.6 46.7
DeepLabv3+ (17arXiv)[DeepLabv3] 72.4 71.6 59.4
SegFormer (21NiPS)[SegFormer] 84.6 85.2 61.3
CloudFU-Net (24TGRS)[CloudFU-Net] 86.9 86.5 62.1
CloudSwinNet (24Energy)[CloudSwinNet] 87.1 87.6 62.3
MFAFNet (25RS)[MFAFNet] 86.7 86.1 61.7
BSANet (25RSA)[BSANet] 86.4 86.5 61.6
FA-CloudSeg (25ISCAIT)[FA-CloudSeg] 88.3 89.2 63.1
@@ IV-D2 Qualitative Comparison
To further demonstrate the superiority of our MPCM-Net in the ground-based cloud image segmentation task, we analyze the results of the comparison methods from a qualitative perspective. We selected some representative samples from the CSRC dataset, including clouds with variant scales, clouds around the sun, and complex boundaries. To comparatively evaluate segmentation performance across cloud scales, Fig. 8 (I) and (II) present sequential frames at 30-second intervals from identical regions, while Fig. 8 (III) demonstrates performance under diverse meteorological conditions and complex boundaries. The first and second columns present the original image and corresponding ground truth masks, respectively, with subsequent columns displaying results from our method and all comparative methods.
@@
Compared to other methods, the proposed MPCM-Net demonstrates superior segmentation performance for ground-based cloud images with variant scales. As shown in the first row of Fig. 8, the segmentation results of U-Net and Deeplabv3+ overlook significant local details due to the highly variable scale distribution of clouds in ground-based cloud imagery, leading to incomplete cloud extraction. While MFAFNet employes parallel dilated convolutions to capture multi-scale cloud features, and CloudFU-Net incorporates an adaptive kernel selection mechanism to adjust receptive field sizes for enhanced multi-scale feature processing, these approaches substantially mitigate the omission of local information in segmentation outputs. Nevertheless, as evident from comparing the first and second rows of Fig. 8, these methods still exhibit incomplete segmentation when processing the same cloud cluster across different frames. In contrast, our MPCM-Net integrates the MPAC, which enhances the extraction of latent features in diverse local regions of multi-scale ground-based cloud imagery. This module strengthens interactions between partial convolution features while adaptively capturing the structural information of multi-scale cloud contexts, thereby achieving optimal segmentation performance. Furthermore, MPCM-Net with MPAC maintains its superior segmentation capability under diverse weather conditions, such as cumulus and cirrus clouds during overcast or rainy periods, as validated by the first row of Fig. 8. Additionally, the third row of Fig. 8 reveals that traditional semantic segmentation algorithms often lose critical boundary details, resulting in subpar cloud segmentation. Furthermore, around the sun, the radiation source causes inaccurate segmentation of surrounding clouds. Addressing the complexity of cloud boundaries, BSANet utilizes a dual-branch structure to simultaneously extract cloud and background features, leveraging complementary background information to enhance boundary feature extraction. MFAFNet employes different attention mechanisms to alleviate the loss of boundary details during decoder upsampling. However, these methods overlook the correlations between local features and global information across different hierarchical levels. Conversely, to deal with the cloud pixel blur caused by sunlight and complex boundaries, the segmentation results of our proposed method ensure the integrity of cloud boundary information, attributed to the M2B incorporating the SSHD. M2B enhances the network’s ability to capture global relationships among local features across layers, compensates for lost contextual information between the encoder and decoder, and improves boundary segmentation in ground-based cloud imagery. In summary, MPCM-Net outperforms existing methods in addressing key challenges in ground-based cloud image segmentation, including scale variation and boundary information loss.
@@ IV-D3 Complexity Comparison
Accurate segmentation of ground-based cloud imagery is essential for capturing critical cloud structure information and constitutes an indispensable component for precise PV power prediction. Furthermore, real-time inference capability is crucial for ultra-short-term PV power forecasting. Consequently, evaluating the accuracy-speed trade-off of different methods in ground-based cloud image segmentation is imperative. To evaluate the computational complexity of our method, we compare parameters, floating-point operations (FLOPs), inference time, and MIoU with relevant approaches on the CSRC dataset. As shown in Table IV, the proposed MPCM-Net achieves the highest performance with the shortest inference time among all evaluated methods. While our approach does not exhibit a parameter advantage compared to the lightweight BSANet, the significant performance gain fully justifies the additional computational overhead. Additionally, the proposed ParAM combines the strengths of linear attention and softmax attention, offering both linear complexity and high-performance feature representation. Consequently, our method yields lower computational costs than self-attention-based segmentation approaches (FA-CloudSeg and CloudSwinNet). Inference time comparisons demonstrate that MPCM-Net achieves near-optimal efficiency, meeting the requirements for ultra-short-term PV power prediction. This efficiency is primarily attributed to the M2B in the decoder, which employs a linear scanning mechanism to enhance global dependencies among local features across network layers while accelerating inference. Therefore, the proposed MPCM-Net establishes an optimal accuracy-speed trade-off for ground-based cloud image segmentation.
TABLE IV: Complexity of Different Comparative Methods on the CSRC Dataset. We Report the Parameters(M), Flops(G), Inference time(ms), and MIoU. $\downarrow$ (or $\uparrow$) indicates lower (or higher) is better. The Best Results are Highlighted in Bold.
@@
U-Net (15MICCAI)[U-Net] 39.5 28.7 26.9 46.7
DeepLabv3+ (17arXiv)[DeepLabv3] 40.4 30.5 27.9 59.4
SegFormer (21NiPS)[SegFormer] 84.7 64.5 42.3 61.3
CloudFU-Net (24TGRS)[CloudFU-Net] 48.8 36.7 28.7 62.1
CloudSwinNet (24Energy)[CloudSwinNet] 68.9 57.6 40.6 62.3
MFAFNet (25RS)[MFAFNet] 34.5 23.1 16.9 61.7
BSANet (25RSA)[BSANet] 4.29 7.5 12.5 61.6
FA-CloudSeg (25ISCAIT)[FA-CloudSeg] 53.6 46.4 36.8 63.1
@@
TABLE V: Comparative analysis of partial attention mechanisms ($\mathrm{P}_{*}$) versus full-channel attention mechanisms ($\mathrm{F}_{*}$) on the CSRC dataset. $\downarrow$ (or $\uparrow$) indicates lower (or higher) is better. The best results are highlighted in bold.
@@
TABLE VI: Ablation experiments of ParCM, ParSM, and ParAM on the CSRC dataset. $\downarrow$ (or $\uparrow$) indicates lower (or higher) is better. The best results are highlighted in bold.

### v2

@@ III-E Loss Function
Ground-based cloud segmentation is intrinsically a binary pixel-level classification task characterized by severe class imbalance. Cloud formations often exhibit significant scale variations and low contrast against the sky background. To accelerate convergence and mitigate bias toward the background class, we employ a compound objective function:
$L_{joint}=\lambda_{1}L_{focal}+\lambda_{2}L_{dice}$ (24)
where $L_{focal}$ denotes the focal loss and $L_{dice}$ denotes the dice loss. $\lambda_{1}$ and $\lambda_{2}$ are both hyperparameters, set to 0.6 and 0.4, respectively.
The Focal Loss addresses the limitations of cross-entropy by introducing a modulating factor that down-weights easy negatives, forcing the model to focus on hard-to-classify boundary pixels. Complementarily, the Dice Loss optimizes the Intersection-over-Union directly, enhancing the model’s sensitivity to the structural integrity of cloud masses.
$\begin{split}L_{focal}=-\frac{1}{M}\sum_{i=1}^{M}(\gamma(1-p_{i})^{\delta}\times p_{i}log(p_{i})+\\ (1-\gamma)p_{i}^{\delta}\times(1-g_{i})log(1-p_{i}))\end{split}$ (25)
$L_{dice}=1-\frac{2\times\sum_{i=1}^{M}p_{i}g_{i}+\varepsilon}{\sum_{i=1}^{M}(p_{i}^{2}+g_{i}^{2})+\varepsilon}$ (26)
where $p_{i}$ and $g_{i}$ denote the prediction and ground truth of sample $i$, respectively. $M$ denotes the number of pixels used in the loss computation, and the index $i$ refers to pixels. $\gamma$ and $\delta$ denote the weight of the positive and negative samples. They are hyperparameters and the values are set to 0.25 and 0.2, respectively, as same as [34]. $\varepsilon$ denotes the smoothing coefficient, which is set to $10^{-5}$ as a hyperparameter to prevent the denominator in equation (26) from being zero.
@@ IV-A Dataset
To bridge this gap and provide the research community with a more challenging and practical benchmark, we developed and publicly released a new dataset CSRC, which constitutes a core contribution of this paper. The solar radiation effect on clouds is a critical factor in the accuracy of PV power forecasting. International cloud classification standards are primarily based on the brightness, color, size, and altitude of clouds. Among these, color variations induced by radiative sources play a significant role in the precision of pixel-level classification in ground-based cloud imagery. In this study, we focus particularly on cloud scale and color attributes, and introduce fine-grained segmentation that incorporates radiative influence, re-categorizing clouds into three classes (see Table I for details). White and gray clouds are common color patterns in the sky, often covering extensive areas, and are represented by white and gray labels, respectively. Pixels within the circumsolar region often appear nearly pure white due to high radiation, with red and blue channel values converging, leading these regions to be classified as overcast. Therefore, it is essential to incorporate fine-grained solar radiation segmentation [24], which is indicated in red. Black cloud systems, typically associated with heavy rain or hail, result in full sky coverage and are detrimental to both ground-based cloud image segmentation and PV power prediction tasks. Such black clouds are ignored from our dataset. Cloud-free sky regions are labeled in blue as background. Table II presents a comparative analysis of the proposed CSRC dataset and existing public benchmarks, evaluating performance across critical dimensions including fine-grained categorization support, radiation source diversity, color attribute availability, spatial resolution, and sample size. Regarding data completeness, it should be noted that the specific sample size is not explicitly documented in Ref[24]; consequently, the designation “N.M.” (Not Mentioned) is adopted in the table to denote this absence.
@@
Dataset Fine-grained Presence of Radiation Source Color Attributes Resolution Sample Size
@@
SWINSEG[42] No No No $\times500$ 115
WAHRSIS[8] No No No $\times600$ 1,013
Fabel et al.[11] No No No $\times512$ 770
Park et al.[35] No No No $\times300$ 1,537
TLCDD[49] No No No $\times512$ 5,000
@@
Shi et al.[39] Yes No No $\times1704$ 300
CSRC (Ours) Yes Yes Yes $\times1260$ 2,330
The images in our CSRC dataset were collected at a meteorological observation station located in Xiqing District, Tianjin, China (geographic coordinates: 117.03° E, 39.10° N). The core acquisition device is an All-Sky Imager (ASI-DC-TK02), which provides a wide-angle field of view exceeding 180° and is housed in a waterproof and dustproof enclosure. The system was configured to automatically capture images at fixed 30-second intervals. All images are stored in RGB color JPG format with a resolution of 1260 × 1260 pixels. The CSRC dataset comprises 2,330 samples; the data were randomly shuffled and divided into training, validation, and test sets in a 7:1:2 ratio. To ensure the fidelity of the ground truth, the annotation process was conducted manually by researchers specializing in cloud analysis utilizing the LabelMe software, under the strict supervision of meteorologists. Representative samples from the CSRC dataset are shown in Fig. 7.
@@ IV-B Evaluation Metrics
As ground-based cloud segmentation fundamentally constitutes a pixel-wise classification task, we employ standard semantic segmentation metrics to quantify model performance. Specifically, we utilize Precision (P), Recall (R), and Mean Intersection over Union (MIoU). All the metrics are derived from true positives (TP), false positives (FP), and false negatives (FN) in experimental results, as specified in the following equations:
@@
where $C$ represents the number of categories in our dataset, and $i=\left\{0,1,2,3\right\}$ represents a total of four categories.
@@ IV-C Implementation Details
The proposed method was implemented in Python 3.9 using PyTorch 1.10.1 on an Ubuntu 21.04 server equipped with an Intel (R) Xeon (R) Gold 5218N CPU @ 2.30 GHz and an NVIDIA GeForce RTX 4090 GPU (24GB VRAM). Training employed an initial learning rate of $10^{-2}$ with the Adam optimizer, configured with momentum 0.9 and weight decay $10^{-3}$ to ensure stable convergence. All input images were uniformly cropped to 512 $\times$ 512 resolution, consistent with the literature [39], to maintain experimental validity. Learning rate decay was applied every 10 epochs, with training termination triggered if no validation loss improvement occurred within 20 consecutive epochs. To enhance model generalization through input diversity, data augmentation strategies including random rotation (0°, 90°, 180°, 270°) and horizontal flipping (probability 0.3) were implemented. The complete training regimen required 8.9 hours for 50 epochs on our dataset.
@@ IV-D Results and Discussions
To rigorously evaluate the performance of MPCM-Net, we conducted extensive comparative experiments against a spectrum of baseline methods, ranging from classical semantic segmentation architectures (U-Net [36], DeepLabv3+ [4], SegFormer [45]) to recent, domain-specific SOTA models (CloudFU-Net [39], CloudSwinNet [38], MFAFNet [13], BSANet [28], and FA-CloudSeg [22]). To ensure a rigorous and fair comparison, all baseline methods were re-implemented and re-trained specifically on the CSRC dataset. We utilized the official source code provided by the original authors to maintain architectural fidelity. We strictly adhered to a unified experimental protocol, ensuring identical data partitioning, augmentation pipelines, and evaluation metrics across all trials. Furthermore, independent hyperparameter tuning was conducted for each comparative method to guarantee their optimal convergence.
@@ IV-D1 Quantitative Comparison
TABLE III: Quantitative Comparison with Different Methods on the CSRC Dataset. $\downarrow$ (or $\uparrow$) indicates lower (or higher) is better. The Best Results are Highlighted in Bold.
@@
U-Net (15MICCAI)[36] 66.7 64.6 46.7
DeepLabv3+ (17arXiv)[4] 72.4 71.6 59.4
SegFormer (21NiPS)[45] 84.6 85.2 61.3
CloudFU-Net (24TGRS)[39] 86.9 86.5 62.1
CloudSwinNet (24Energy)[38] 87.1 87.6 62.3
MFAFNet (25RS)[13] 86.7 86.1 61.7
BSANet (25RSA)[28] 86.4 86.5 61.6
FA-CloudSeg (25ISCAIT)[22] 88.3 89.2 63.1
@@ IV-D2 Qualitative Comparison
Qualitative comparisons, presented in Fig. 8, provide visual corroboration of the quantitative metrics. We selected challenging samples featuring multi-scale cloud variations, circumsolar occlusion, and complex boundaries. To comparatively evaluate segmentation performance across cloud scales, Fig. 8 (I) and (II) present sequential frames at 30-second intervals from identical regions, while Fig. 8 (III) demonstrates performance under diverse meteorological conditions and complex boundaries.
@@
As shown in the first and second rows of Fig. 8, cloud formations exhibit significant scale variations across sequential frames. Classical methods (U-Net, DeepLabv3+) frequently omit fine-grained local details of smaller cloud fragments. While CloudFU-Net utilizes adaptive kernel selection to adjust receptive fields, it still exhibits temporal inconsistency when segmenting the same cloud cluster across 30-second intervals. MPCM-Net, leveraging the MPAC module, demonstrates superior consistency, accurately resolving both large-scale stratiform clouds and minute cumulus fragments. The third row of Fig. 8 highlights the challenge of circumsolar regions, where high radiation intensity blurs the spectral boundary between cloud and sky. Traditional algorithms often fail here, misclassifying the aureole as cloud. BSANet and MFAFNet attempt to ameliorate this via dual-branch structures and boundary-attention mechanisms, respectively. However, they occasionally overlook the correlation between local boundary features and the global semantic context. The M2B block with SSHD enhances the network’s capability to capture global dependencies among local features, effectively ”contextualizing” the ambiguous boundaries and resulting in sharper, more accurate segmentation maps even in saturated circumsolar zones.
@@ IV-D3 Complexity Comparison
For practical deployment in ultra-short-term PV power forecasting, inference speed is as critical as accuracy. Table IV compares the parameters, Floating Point Operations (FLOPs), and inference time of all methods. MPCM-Net establishes an optimal accuracy-efficiency trade-off. It achieves the highest segmentation accuracy while maintaining the lowest inference latency. While our approach does not exhibit a parameter advantage compared to the lightweight BSANet, the significant performance gain fully justifies the additional computational overhead. Notably, it incurs significantly lower computational costs than Transformer-based approaches (FA-CloudSeg, CloudSwinNet). This efficiency is a direct consequence of the M2B architecture’s linear scaling law, which avoids the heavy computational burden of the query-key-value calculations inherent in SA mechanisms. Consequently, MPCM-Net is uniquely positioned to meet the real-time requirements of high-frequency ground-based cloud observation systems.
TABLE IV: Complexity of Different Comparative Methods on the CSRC Dataset. We Report the Parameters(M), Flops(G), Inference time(ms), and MIoU.
@@
U-Net (15MICCAI)[36] 39.5 28.7 26.9 46.7
DeepLabv3+ (17arXiv)[4] 40.4 30.5 27.9 59.4
SegFormer (21NiPS)[45] 84.7 64.5 42.3 61.3
CloudFU-Net (24TGRS)[39] 48.8 36.7 28.7 62.1
CloudSwinNet (24Energy)[38] 68.9 57.6 40.6 62.3
MFAFNet (25RS)[13] 34.5 23.1 16.9 61.7
BSANet (25RSA)[28] 4.29 7.5 12.5 61.6
FA-CloudSeg (25ISCAIT)[22] 53.6 46.4 36.8 63.1
@@
TABLE V: Comparative analysis of partial attention mechanisms ($\mathrm{P}_{*}$) versus full-channel attention mechanisms ($\mathrm{F}_{*}$) on the CSRC dataset.
@@
TABLE VI: Ablation experiments of ParCM, ParSM, and ParAM on the CSRC dataset.